# Predictable Drifts in Collective Cultural Attention: Evidence from Nation-Level Library Takeout Data

### v2

@@ Abstract
Predicting changes in consumer attention for cultural products, such as books, movies, and songs, is notoriously difficult. Past research on predicting the popularity of individual products suggests the existence of intrinsic prediction limits. However, little is known about the limits for predicting collective attention across cultural products. Here, we analyze four years of nationwide library loan data for approximately 2 million individuals, comprising over 100 million loans of more than 660,000 unique books. We find that culture, as measured by popularity distributions of loaned books, drifts continually from month to month at a near-constant rate, leading to a growing divergence over time, and that drifts vary between different book genres. By linking book loans to registry data, we investigate the influence of age, sex, educational level, and geographical area on cultural drift, finding heterogeneous effects from the different demographic groups. Our findings have important implications for market forecasting and developing robust recommender systems, highlighting the need to account for specific drift dynamics for different types of items and demographic groups.
@@
cultural markets, predictability, library data, social systems, data drift
Cultural products drive billion-dollar industries, influence public discourse, and shape our personal and collective identities. Yet, their popularity remains notoriously hard to predict, even despite advances in predictive modeling and the growing availability of large-scale cultural consumption data [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. This unpredictability coincides with extremely unequal distributions of popularity, where a small number of products capture a large share of consumer attention [12, 13, 14, 15, 16]. Seminal work suggests that the unpredictability and inequality observed in cultural markets can be explained by high levels of social influence [17, 18]. Strong social influence on individual decisions can produce information cascades in cultural markets and other social systems, where large numbers of consumers simultaneously attend to the same cultural items [19, 20, 21]. Multiple studies have systematically analyzed the dynamics of collective attention, revealing characteristic patterns in how popular products and major events enter into and fade from collective attention [22, 23, 24, 25, 26, 27].
@@
However, existing approaches have not been able to address important questions about collective consumer attention in cultural markets. How much does collective attention change on aggregate? Are these changes universal across products and consumer groups? Are they predictable? Answers to these questions have implications for our theoretical understanding of cultural markets and for practical applications involving models of market dynamics, for instance, in developing robust recommender systems. Resolving them requires comprehensive cultural consumption data that moves beyond the current focus on hit products as well as an analytical framework for quantifying changes in collective attention at the market level.
@@
Here, we use a unique nation-scale book loan dataset which contains records from all public libraries in Denmark, featuring 100 million book loans from 2 million individuals (over one third of the Danish population), to study how cultural attention evolves at a societal scale. In Denmark, library borrowing has a low barrier to entry since library services are well-funded, free to use, and accessible anywhere in the country. In addition, Danish public libraries have a wide coverage, and any material present in the library system can be ordered from any local library. For these reasons, libraries serve as a primary means of accessing books for Danish citizens. According to the best available data on book sales in Denmark, the number of book loans from public libraries is consistently almost three times the number of book sales (see SI Fig. LABEL:book_sales_loans_comparison based on [28]). Further, Denmark is one of the most digitalized countries in the world, making it possible to link each individual book loan to registry data, which provides information about library users’ demographic characteristics, such as age, sex, education level, zip code, etc. Given the unprecedented scope of this dataset, we can identify changes in collective consumer attention for a wide range of books across a large number of consumers.
@@
Our approach differs fundamentally from prior work on collective attention. Rather than focusing on the individual trajectories of popular items in isolation, we analyze the changes in consumer attention across all items in the market, which we refer to as ‘drifts’ in collective attention. This novel approach to measuring changes in collective attention eliminates the need to filter away less popular products, capturing attention dynamics across the full popularity spectrum. In addition, it requires no assumptions about how individual items gain or lose consumer attention, accommodating diversity in popularity trajectories.
@@
Our analysis is split up into three parts. First, we quantify drift in collective attention, finding that culture has a constant turnover rate across months. Second, we dig into which demographic groups contribute to this drift, finding that age, sex, and place of residence have a large influence. Third, we account for the contributions of each unique book title to collective attention drift. We use our findings to build a simple model that can accurately estimate drift without needing to account for individual product trajectories.
1 Results
We use comprehensive records of all Danish public library loans (excluding university and school libraries) from start 2020 to end 2023. Each entry in our dataset corresponds to the loan of a single book and contains information on the date of the loan, a unique identifier of the book, a pseudonymized loaner ID, the library where the book was loaned, as well as metadata on the books and loaners (see Fig. 1A). The dataset contains more than 100 million unique book loans from 513 libraries, including eBooks from digital services (see SI Fig. LABEL:material_type_distribution for the distribution of material types) and features loans from over 45% of the entire Danish population. A single book title has multiple identifiers for different media (physical book, audiobook, etc.) and editions. Throughout our analysis, we treat each variant of a book as a single entity (see Methods for details on how we identify individual books). We link loaner IDs to national registry data to obtain sociodemographic information, including age, sex, education level, and area of residence for each loaner (see Methods).
@@ 1 Results
Figure 1: Dataset and measurement overview. A, Each entry in the dataset corresponds to the loan of a single book, illustrated as vertical lines. Some loaner attributes (sex and birthdate) are static, whereas others (loaner residence and education level) are dynamic. We aggregate loans into monthly time periods and measure the local (between successive periods) and global (between a given period and a reference period) drifts in cultural attention. B, Aggregate daily number of loans. A nationwide COVID-19 lockdown, effective March 13, 2020, and lasting until May 18, 2020, shut down all libraries. The few loans of physical books during this period can be attributed to some libraries putting out ‘book bags’ for their loaners. A partial lockdown, initiated on December 9, 2020, limited public access to library interiors in 38 municipalities. During these lockdown periods, there was a noticeable decrease in the number of books borrowed from physical libraries and an increase in digital loans. The inset shows an example of the weekly fluctuations in takeout numbers over two weeks from 24/01/2022 to 06/02/2022. C, Weekly number of loans for 20 popular books in the dataset. D, Complementary cumulative distribution function (CCDF) of the number of book loans for the more than 660,000 books over the entire four-year period. The inset shows the CCDFs for three main categories of books. E, Number of loans for different sociodemographic subgroups.
@@
To quantify drifts in library book consumption, we use the Jensen-Shannon Divergence (JSD). The JSD is an information-theoretic measure that is well-suited to measure changes in sequences of frequency distributions [29]. In the social science literature, JSD and related measures have mainly been used to analyze text sequences [30, 31, 32, 33, 34, 35, 36, 37]. In this study, we use it on sequences of monthly book popularity distributions (see Methods for technical details). We use months as the aggregation level since it allows us to capture changes in consumption without being overly sensitive to minor variations in book popularity that do not reflect meaningful changes in consumer attention. If the JSD between loans in two months is 0, it means the popularity distributions are identical; if it is 1, entirely different books have been loaned in the two months. Values between 0 and 1 represent a partial reallocation of collective attention.
@@
We distinguish between two main types of drift: local drift, which quantifies the JSD between a given month and the preceding month, and global drift, which measures drifts between a reference month and all other months (Fig. 1A). The two drift types reveal distinct patterns in how library book popularity evolves. While local drifts capture immediate shifts in book consumption, global drifts capture the longer-term changes to the loaning behavior.
@@
Over the 4-year-span of our dataset, we observe variation in the number of library loans (Fig. 1B). In particular, two COVID-19-related lockdowns restricted library access in Denmark, resulting in a dramatic drop in how many physical books were loaned as well as an increase in digital book loans. Additionally, citizens tend to loan more books in the Summer holidays and less when libraries are typically closed (on Sundays and in the Winter holidays). To minimize noise and the effects of COVID-19, our analysis focuses on the period after the last lockdown (from May 2021 to December 2023). Individual book trajectories reveal considerable heterogeneity in how popularity evolves over time (Fig. 1C). While some books experience sharp peaks followed by steady decline, others maintain more stable popularity levels or exhibit complex patterns, including delayed and recurring peaks in popularity.
@@
Collective consumer attention is highly unequally distributed among the 660,000 books in our dataset (Fig. 1D). An unequal popularity distribution is not unique to our dataset but is a commonly observed pattern across cultural markets [12, 13, 14, 15, 16]. In our data, most books are loaned less than 10 times, while only around 1,000 books attract more than 10,000 loans. Looking across book genres, loans appear to be most unequally distributed among adult fiction, and most equally distributed for nonfiction books (see inset in Fig. 1D). The shape of the popularity distributions is constant across months (see SI Fig. LABEL:book_popularity_distributions_months). This means that drifts in collective attention between two months capture changes in which books hold which positions in the rank-frequency distribution rather than simply changes in the shape of the distribution, such as increased or decreased diversity in book loans.
@@
Loaner subgroups are not equally represented in this dataset (Fig. 1E). For instance, there is a clear gap between the sexes, with female loaners lending more books. Additionally, book types are not evenly distributed among loaner subgroups, especially among different age segments. Unsurprisingly, young loaners (0-17 years) almost exclusively loan children’s books. Loaners aged 30 to 45 take out many children’s books and relatively few adult fiction books, whereas 65+ loaners borrow very few children’s books and many adult fiction books. In SI Fig. LABEL:book_popularity_distributions_loaner_groups, we show the distribution of loans across different book types for each subgroup. Further, SI Fig. LABEL:book_popularity_distributions_loaner_groups_months demonstrates that these distributions are stable over time.
1.1 Patterns of collective drifts
@@
We start our analysis at the most general level by quantifying drifts in popularity distributions across all books and loaners. Later, we will segment our analysis and investigate the contribution of different book categories and population demographics to drift. Fig. 2A shows the temporal dynamics of local and global drifts. Local drifts are approximately constant between successive months, punctuated only by small recurring peaks from November to January (red, solid line Fig. 2A). These deviations stem from season-specific books (predominantly Christmas-themed books) becoming popular in November and December. Global drifts, however, grow continually from the reference month of May 2021 (blue, solid line Fig. 2A), and show similar season-specific bumps. Taken together, this suggests two primary mechanisms of drift: a gradual, persistent shift in loaner preferences and a seasonal component causing momentary increases in collective attention.
@@
Figure 2: Local and global drifts in cultural attention. A, Local (red) and global (blue) drifts across all book categories and loaner demographics. We estimate JSD in bits to ensure that drifts are bounded between 0 and 1. B-I, Local and global drifts split up according to population demographics and book category. To minimize the effects of random fluctuations and noise, we focus only on the top 10,000 most popular books. B-E, shows drifts for the top 10,000 most popular books, while panels F-I show drifts for the top 10,000 most popular adult books. Shaded areas around lines (barely visible) show $2\sigma$ bootstrap errors.
@@
To explore whether patterns are similar across different book categories, we compute local and global drifts for three broad categories of books, which are available in our dataset: adult fiction, adult nonfiction, and children’s books. For all book categories, local drifts are stable, meaning that collective attention has a constant rate of drift, while global drifts grow over time (dashed and dotted lines Fig. 2A). However, drift patterns differ in notable ways. Specifically, adult nonfiction books have a larger rate of local drift compared to adult fiction and children’s books. One explanation for this is that a large proportion of nonfiction books are rarely loaned, as seen in Fig 1C. When we exclude infrequently loaned books, the local drift rate for nonfiction books becomes less extreme, but it is still higher than drifts in fiction and children’s books (see SI Fig. LABEL:drifts_top10k). For global drift, we observe different dynamics. Drift in children’s books grows at a considerably slower rate than in adult books. Lastly, for both local and global drifts, the seasonal peaks (around Christmas) are most pronounced for children’s books, less so for adult nonfiction books, and nearly absent for adult fiction books. In the SI, we show that these patterns are qualitatively robust: across different pairs of months (SI Fig. LABEL:drifts_heatmap_all_books-LABEL:drifts_heatmap_children), using generalized measures of drift (see SI Sec 1.1 and SI Figs. LABEL:drifts_alternative_measures_all_books-LABEL:drifts_alternative_measures_children), and for different temporal aggregation levels (SI Figs. LABEL:drifts_week-LABEL:drifts_quarter_top10k). We further show that patterns are qualitatively robust, although drift magnitudes are larger, across COVID lockdowns when we include data from January 2020 (SI Figs. LABEL:drifts_monthly_from-2020-LABEL:drifts_across-covid_comparison).
1.2 Drifts split up according to population subgroups
@@
So far, we have investigated drifts across the entire Danish population. Here we investigate how dynamics vary across demographic subgroups. Prior work has shown that sociodemographic groups, defined by sex, age, education level, and area of residence to exhibit differences in the diversity of cultural consumption [38, 39]. Yet, to our knowledge, no study has quantified how the collective consumption of these segments evolves over time. To compare different subgroups of the loaner population, we account for the fact that drifts computed from raw frequencies capture both changes to the underlying collective attention and random fluctuations in the popularity of less popular books. To eliminate the effects of random fluctuations, we compute drifts exclusively for the top 10,000 most popular books within each sociodemographic subgroup and correct for any residual bias in drift estimation using bootstrap methods [40] (see Methods for details). For the comparison of loaners in different age bins, we exclude loaners below the age of 30 as loaners from 0 to 17 almost exclusively lend children’s books, and loaners between 18 and 29 have very few loans.
@@
We find substantial differences in drifts between loaners of different ages (Fig. 2B). Both local and global drifts are considerably lower for 30-45-year-old loaners than for older ones, and the seasonal component is stronger. For sex (Fig. 2C), we find females have a larger global drift and stronger recurring seasonal peaks, while differences in local drift are minuscule. Additionally, it is worth noting that the frequency of seasonal peaks differs between males and females. While females show seasonal peaks centered around Christmas, males, in addition, also have summer peaks. For other sociodemographic subdivisions (education level and area of residence), differences in local drifts are not pronounced, and we observe only small to moderate differences between global drifts over time (Figs. 2D-E). Interestingly, seasonally recurring peaks are more pronounced for 30-45-year-olds and females, this is partially due to these groups loaning more children’s books (see Fig. 1E).
@@
To investigate the role of children’s books in drift for individual demographic groups, we compare the results from Fig. 2B-E with drifts specifically calculated for the 10,000 most popular adult books (i.e., disregarding loans of children’s books). Here, we observe that the differences in global drifts between age groups are smaller when we only consider drifts in adult books (Fig. 2F). In addition, disregarding children’s books makes the differences in global drift more prominent between sex-, education-, and residence area-groups (Fig. 2G-I). Here we observe that drift is higher for females, for more educated individuals (Bachelor or higher), and for individuals living in large cities. These differences are robust when books are further split up into adult fiction and adult nonfiction (SI Fig. LABEL:drifts_subgroups_adult_fic_nonfic). When it comes to seasonal peaks, we find that removing children’s books eliminates the regularly occurring peaks for 30-45-year-olds (Fig. 2F), and both sexes (Fig. 2G). This is the case for both global and local drifts, and aligns well with younger demographics being more likely to have children.
1.3 Contributions from specific books
@@
Until now, we have considered books collectively; here, we untangle the importance of individual items on drift and study what role individual books play in collective attention drift. We do this as much of the focus in both media and previous studies [1, 2, 3, 4, 8, 9, 25, 6] is on big hits that occupy our collective attention at any given moment; however, it is unclear how important individual items are for explaining overall drifts in collective attention. To investigate this, we measure how much each book contributes to (nation-scale) local and global drifts. For each month, we divide all books into five broad groups defined by the size of their contributions (see Methods for details). I.e., the more popular a book is, and the more its relative popularity fluctuates between months, the higher its contribution (see Methods). Groups are non-overlapping and defined as: i) the top 100 most contributing books, ii) the top 101 to 1000 most contributing books, iii) top 1001 to 10.000, iv) top 10.001 to 50.000, and v) books beyond top 50.000. Using these groups, we explore how many books we would need to account for a given proportion of the drift over time and how consistent the contributions of individual books are.
We start by examining how much each group of books contributes to the drift in collective attention over time (Fig. 3). Surprisingly, we find that the 100 books with the most significant shifts in their popularity consistently account for only around 10% of both local and global drifts (Fig. 3A-B). As such, even if it were possible to identify these books and predict changes in their popularity, 90% of the total shift in collective attention would remain unexplained, unless less influential books were also included. For local drift, the top 10,000 books that contribute most to drift account for less than 50% of the total change in behavior. This shows the importance of considering the full ‘cultural system’ and not just individual product trajectories when analyzing patterns of consumer attention. When it comes to seasonal variation in local drift (Fig. 3A) we observe increases in the relative contributions of the top 101-1,000 most increases, but the top 100 books contribute less to this seasonal pattern. For global drift, we find that the relative contributions of the top 1,000 books decline over time (Fig. 3B). In other words, the 1,000 books with the largest changes in popularity are less important in explaining collective attention drifts for months, the further we move from the reference month. Instead, shifts in the popularity of especially the top 1,001-10K most contributing books account for a larger proportion of global drifts over time. Nonetheless, even when accounting for the top 10.000 most influential books, we explain less than 70% of global drift over 2.5 years. This illustrates that books in the long tail of the popularity distribution, in a collective fashion, are impactful in driving changes in collective attention. In other words, when analyzing the cultural popularity of items, it is vital to consider books across all contribution groups.
@@
Figure 3: Contributions of individual books to drift dynamics. A-B, Distributions of relative contributions of individual books to local drift (A) and global drift (B) over time. C, Probability that a book with a given level of contribution to the local drift in a month will have a certain level of contribution to the local drift in the next month. Probabilities are averaged across all pairs of months. D, Monthly popularity for the 1,000 books with the highest contributions to the global drift in October 2023, sorted by the month where their popularity peaks.

### v3

@@ Abstract
Predicting changes in consumer attention for cultural products, such as books, movies, and songs, is notoriously difficult. Past research suggests intrinsic limits for predicting consumer attention towards individual products. However, little is known about the limits for predicting shifts in collective attention. Here, we analyze five years of nationwide library loan data for almost 3 million individuals, comprising over 136 million loans of more than 750,000 unique titles. We find that culture, as measured by popularity distributions of loaned books, drifts continually from month to month at a near-constant rate, leading to a growing divergence over time, and that drift varies between book genres. By linking book loans to registry data, we investigate the influence of age, sex, educational level, and residential area type on cultural drift, finding heterogeneous effects. Our findings have important implications for market forecasting and algorithmic recommender systems, highlighting the need to account for drift dynamics.
@@
cultural markets, predictability, human behavior, social systems, data drift
Cultural products drive billion-dollar industries, influence public discourse, and shape our personal and collective identities. Predicting which products will capture the attention of audiences holds significant economic and cultural value. Yet, it is notoriously difficult to obtain accurate predictions, even despite advances in predictive modeling and the growing availability of large-scale cultural consumption data [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Seminal work suggests that this difficulty reflects not only data or modeling limitations, but more fundamental limits to prediction in cultural markets, where an abundance of products compete for scarce attention, and consumer choices are strongly shaped by social influence, making outcomes highly contingent and path-dependent [12, 13]. These features are not unique to cultural markets but apply to other domains like news and online content where consumption and attention are meaningfully collective phenomena. Across these settings, a small number of products typically captures a disproportionately large share of consumer attention [14, 15, 16, 17, 18]. Most studies of collective attention focus on the most popular products and major events that capture a large share of our collective attention [19, 20, 21, 22, 23, 24].
@@
Existing research offers little insight into how collective attention changes across an entire market. Understanding these system-level dynamics is important because they may reveal regularities and constraints that remain unobserved when focusing only on products in isolation. Several important questions remain open. How fast does collective attention change overall? Are changes to collective attention universal across products and consumer groups? Are they predictable? Answers to these questions have implications for our theoretical understanding of cultural market dynamics and for practical applications involving models of market dynamics, such as developing robust recommender systems. Resolving them requires comprehensive cultural consumption data that moves beyond the current focus on hit products as well as an analytical framework for quantifying changes in collective attention at the market level.
Here, we use a unique nation-scale book loan dataset which contains records from all public libraries in Denmark, featuring 136 million book loans from 3 million individuals (almost half of the Danish population), to study how cultural attention evolves at a societal scale. While libraries are not a traditional economic market per se, as individuals do not have to pay to loan books, books in a library compete in what is often called an ‘attention market’. In these markets human attention is the limited commodity which these books compete for. We study how attention changes in this system over time, and which factors drive the change.
@@
In Denmark, library borrowing has a low barrier to entry since library services are well-funded, free to use, and accessible anywhere in the country. In addition, Danish public libraries have a wide coverage, and any material present in the library system can be ordered from any local library. For these reasons, libraries serve as a primary means of accessing books for Danish citizens. According to the best available data on book sales in Denmark, the number of book loans from public libraries is consistently almost three times the number of book sales (see Supplementary Fig. S8 based on [25]). In addition, past work has established that loans in our data accurately reflect loaners’ self-reported literary tastes, suggesting that library loans is a good indicator of broader book consumption [26]. Further, Denmark is one of the most digitalized countries in the world, making it possible to link each individual book loan to registry data, which provides information about library users’ demographic characteristics, such as age, sex, education level, zip code, etc. We use the unprecedented scope of this dataset to identify changes in collective consumer attention for a wide range of books across a large number of consumers.
@@
Our approach differs fundamentally from prior work on collective attention. Rather than focusing on the individual trajectories of popular items in isolation, we analyze the changes in consumer attention across all items in the market, which we refer to as drift in collective attention. We borrow the term ‘drift’ from the machine learning literature to describe the shifts in the underlying data distributions that machine learning model are trained on, which can negatively affect model performance [27, 28]. Our approach to measuring changes in collective attention eliminates the need to filter away less popular products, capturing attention dynamics across the full popularity spectrum. In addition, it requires no assumptions about how individual items gain or lose consumer attention, accommodating diversity in popularity trajectories.
@@
Our analysis is split up into three parts. First, we quantify drift in collective attention, finding that culture has a constant turnover rate across months with seasonal variation. Second, we dig into which demographic groups contribute to this drift, finding that age and sex have a large influence. Third, we account for the contributions of each unique book title to collective attention drift. We use our findings to build a simple model that can accurately estimate the future magnitude of drift without needing to account for individual product trajectories.
Results
@@
Figure 1: Dataset and measurement overview. A, Each entry in the dataset corresponds to the loan of a single book, illustrated as vertical lines. Some loaner attributes (sex and birthdate) are static, whereas others (loaner residence and education level) are dynamic. We aggregate loans into monthly time periods and measure the local (between successive months) and global (between a given period and a reference month) drift in cultural attention. B, Aggregate daily number of loans. A nationwide COVID-19 lockdown, effective March 13, 2020, and lasting until May 18, 2020, shut down all libraries. The few loans of physical books during this period can be attributed to some libraries putting out ‘book bags’ for their loaners. A partial lockdown, initiated on December 9, 2020, limited public access to libraries in 38 municipalities. During these lockdown periods, there was a noticeable decrease in the number of books borrowed from physical libraries and an increase in digital loans. The inset shows an example of the weekly fluctuations in takeout numbers over two weeks from 24/01/2022 to 06/02/2022. C, Weekly number of loans for 20 popular books in the dataset. The y-axis is discontinued because of one book with a very high number of loans around Feb, 2024. D, Complementary cumulative distribution function (CCDF) of the number of book loans for the more than 750,000 books over the entire five-year period. The inset shows the CCDFs for three main categories of books. E, Number of loans for different sociodemographic subgroups.
We use comprehensive records of all Danish public library loans (excluding university and school libraries) for five years from start 2020 to end 2024. Each entry in our dataset corresponds to the loan of a single book and contains information on the date of the loan, a unique identifier of the book, a pseudonymized loaner ID, the library where the book was loaned, as well as metadata on the books and loaners (see Fig. 1A and Methods). The dataset contains more than 136 million unique book loans from 513 libraries, including eBooks from digital services (see Supplementary Fig. S9 for the distribution of loans across material types) and features loans from almost 50% of the Danish population. Our analysis focuses on three broad book genre categories: general audience fiction, general audience nonfiction, and children’s books, which encompasses both children’s and young adult literature. A single book title has multiple identifiers for different media (physical book, audiobook, etc.) and editions. For our main results, we treat each variant of a book as a single entity (see Methods for details on how we identify individual titles). We link loaner IDs to national registry data, in a privacy-preserving manner, to obtain sociodemographic information, including age, sex, education level, and area of residence for each loaner (see Methods).
@@
To quantify collective attention drift in library book consumption, we use the Jensen-Shannon Divergence (JSD). The JSD is an information-theoretic measure that is well-suited to measure changes in sequences of frequency distributions [29]. Within the social science literature, JSD and related measures have mainly been used to analyze changes in text [30, 31, 32, 33, 34, 35, 36, 37]. In this study, we use it on sequences of monthly book popularity distributions (see Methods for technical details). We use months as the aggregation level in our main results since it allows us to capture changes in consumption without being overly sensitive to minor variations in book popularity that do not reflect meaningful changes in consumer attention. A JSD of 0 indicates the popularity distributions are identical between loans in two months; a JSD of 1, means entirely different books have been loaned. Values between 0 and 1 represent a partial reallocation of collective attention.
@@
We distinguish between two main types of drift: local drift, which quantifies the JSD between a given month and the preceding month, and global drift, which measures drift between a reference month and all subsequent months (Fig. 1A). The two drift types reveal distinct patterns in how library book popularity evolves. While local drift captures immediate shifts in book consumption, global drift quantifies long-term changes to loaning behavior.
@@
Over the 5-year-span of our dataset, we observe variation in the number of library loans (Fig. 1B). In particular, two COVID-19-related lockdowns restricted library access in Denmark, resulting in a dramatic drop in how many physical books were loaned as well as an increase in digital book loans. Additionally, citizens tend to loan more books in the Summer holidays and less when libraries are typically closed (on Sundays and in the Winter holidays). To minimize noise and the effects of COVID-19, our analysis focuses on the period after the last lockdown (from May 2021 to December 2024). There is considerable heterogeneity in how the popularity of individual titles evolves over time, reflecting the difficulty of predicting individual trajectories (Fig. 1C). While some books experience sharp peaks followed by steady decline, others maintain more stable popularity levels or exhibit complex patterns, including delayed and recurring peaks in popularity.
@@
Collective consumer attention is highly unequally distributed among the over 750,000 titles in our dataset (Fig. 1D). An unequal popularity distribution is not unique to our dataset but is a commonly observed pattern across cultural markets [14, 15, 16, 17, 18]. Loans appear to be most unequally distributed among fiction books targeted a general audience, and most equally distributed for nonfiction books (see inset in Fig. 1D). The shape of the popularity distributions is constant across months (see Supplementary Fig. S10). This means that drift in collective attention between two months capture changes in which books hold which positions in the rank-frequency distribution rather than simply changes in the shape of the distribution, such as increased or decreased diversity in book loans.
@@
Loaner subgroups are not equally represented in this dataset (Fig. 1E). For instance, there is a clear gap between the sexes, with women lending more books. Additionally, book types are not evenly distributed among loaner subgroups, especially among different age segments. Unsurprisingly, young loaners (0-17 years) almost exclusively loan children’s books. Loaners aged 30 to 45 take out many children’s books and relatively few general audience fiction books, whereas 65+ loaners borrow very few children’s books and many general audience fiction books. In Supplementary Fig. S11, we show the distribution of loans across different book types for each subgroup. Further, Supplementary Fig. S2 demonstrates that these distributions are stable over time.
Patterns of collective drift
@@
We start our analysis at the most general level by quantifying drift in popularity distributions across all books and loaners. Fig. 2A shows the temporal dynamics of local and global drift. Local drift is characterized by being near-constant between successive months, punctuated only by small recurring peaks during Winter periods (November to January; red, solid line Fig. 2A). These deviations stem from season-specific titles, which are predominantly Christmas-themed, becoming popular in November and December (see Supplementary Section 1.3). Global drift, however, grows continually from the reference month of May 2021 (blue, solid line Fig. 2A), and show similar season-specific bumps. Taken together, this suggests two primary mechanisms of drift: a gradual, persistent shift in loaner preferences and a seasonal component causing momentary increases in collective attention.
@@
Figure 2: Local and global drift in cultural attention. A, Local (red) and global (blue) drift across all book categories and loaner demographics. We estimate JSD in bits to ensure that drift is bounded between 0 and 1. To pinpoint the contribution to drift from different book categories we use partially overlapping categories. ‘All books’ contains both ‘general audience’ and ‘children’s’ categories, and ‘general audience’ (GA) is further split into ‘fiction’ and ‘nonfiction’ categories. B-I, Local and global drift split up according to population demographics and book category. To minimize the effects of random fluctuations and noise, we focus only on the top 10,000 most popular books. B-E, shows drift for the top 10,000 most popular books, while panels F-I show drift for the top 10,000 most popular general audience books. Shaded areas around lines (barely visible) show $2\sigma$ bootstrap errors.
@@
To explore whether patterns are similar across different book categories, we compute local and global drift for three broad categories of books, which are available in our dataset: general audience fiction, general audience nonfiction, and children’s books. For all book categories, local drifts are stable, meaning that collective attention has a constant rate of drift, while global drift grows over time (dashed and dotted lines in Fig. 2A). However, drift patterns differ in notable ways. Specifically, general audience nonfiction books have a larger rate of local drift compared to general audience fiction and children’s books. One explanation for this is that a large proportion of nonfiction books are rarely loaned, as seen in Fig 1D, inset. When we exclude infrequently loaned books, the local drift rate for nonfiction books becomes less extreme, but it is still higher than drift for fiction and children’s books (see Supplementary Fig. S13). For global drift, we observe different dynamics. Drift in children’s books grows at a considerably slower rate than in general audience books. Lastly, for both local and global drift, the seasonal peaks (around Christmas) are more pronounced for children’s books, less so for general audience nonfiction books, and nearly absent for general audience fiction books.
We assess the robustness of our findings through a series of complementary analyses. The observed patterns are qualitatively robust when using generalized measures of drift (Supplementary Section 1.1 and Supplementary Figs. S14-S19) and for different temporal aggregation levels other than months (Supplementary Figs. S19-S22). Our findings also differ substantially from the dynamics produced by a null model in which changes arise solely from random fluctuations in the popularity of books (Supplementary Section 1.2). Another potential concern is that our results could be affected by the limited supply of books in libraries, as we quantify drift from the number of loans rather than the number of requests for a book. Indeed digital books (e-books and audiobooks), which have an unlimited supply, typically have a high number of loans for a shorter time than physical books (Supplementary Fig. 23), suggesting that the supply of physical books cannot always follow the demand. However, drift patterns measured separately for physical and digital books are nearly identical (Supplementary Fig. 24), indicating that supply constraints do not drive our results.
Drift split up according to population subgroups
@@
Prior work has shown that sociodemographic groups, defined by sex, age, education level, and area of residence to exhibit differences in the diversity of cultural consumption [38, 39]. Yet, to our knowledge, no study has quantified how the collective attention of these segments evolves over time. To compare different subgroups of the loaner population, we account for the fact that drift computed from raw frequencies capture both changes to the underlying collective attention and random fluctuations in the popularity of less popular books. To eliminate the effects of random fluctuations, we compute drift exclusively for the top 10,000 most popular books within each sociodemographic subgroup and correct for any residual bias using bootstrap methods [40] (see Methods for details). For the comparison of loaners in different age bins, we exclude loaners below the age of 30 as loaners from 0 to 17 almost exclusively lend children’s books, and loaners between 18 and 29 have very few loans.
@@
We find substantial differences in drift between men and women, where the latter have a larger global drift and stronger recurring seasonal peaks, while differences in local drift are minuscule (Fig. 2B). Additionally, it is worth noting that the frequency of seasonal peaks differs between men and women. While women show seasonal peaks centered around Christmas, men, in addition, also have summer peaks. For loaners of different ages we also find large differences in drift (Fig. 2C). Both local and global drift are considerably lower for 30-45-year-old loaners than for older ones, and the seasonal component is stronger. For other sociodemographic subdivisions (education level and area of residence), differences in local drift are not pronounced, and we observe only small to moderate differences between global drift over time (Figs. 2D-E).
@@
To investigate the role of children’s books on drift for individual demographic groups, we compare the results from Fig. 2B-E with drift specifically calculated for the 10,000 most popular general audience books (i.e., disregarding loans of children’s books). Here, we observe that the differences in global drift between men and women becomes larger when we only consider drift in general audience books (Fig. 2F). In addition, disregarding children’s books also makes the differences in global drift more prominent for education- and residence area-groups (Fig. 2H-I). Here we observe that drift is higher for women, for more educated individuals (Bachelor or higher), and for individuals living in large cities. Differences between age groups, however, get smaller (Fig. 2G). Disregarding children’s books reduces drift to the same order of magnitude as sex, where differences between older demographics (65+ year-olds) and younger demographics (30-45-year-olds) become similar to that of men and women. These differences are robust when books are further split up into general audience fiction and general audience nonfiction (Supplementary Fig. S25). When it comes to seasonal peaks, we find that removing children’s books eliminates the regularly occurring peaks for both men and women (Fig. 2F), and 30-45-year-olds (Fig. 2G), consistent with our previous finding that children’s books drive the seasonal peaks nationally.
Contributions from individual books
@@
Until now, we have considered books collectively; here, we untangle the importance of individual items on drift and study what role individual books play in collective attention drift. We do this as much of the focus in both media and previous studies is on big hits that occupy our collective attention at any given moment [1, 2, 3, 4, 8, 9, 22, 6]. However, it is unclear how important individual items are for explaining overall drift in collective attention. To investigate this, we measure how much each book contributes to (nation-scale) local and global drift. For each month, we divide all books into five broad groups defined by the size of their contributions. I.e., the more popular a book is, and the more its relative popularity fluctuates between months, the higher its contribution (see Methods for details). Groups are non-overlapping and defined as: i) the top 100 most contributing books, ii) the top 101 to 1000 most contributing books, iii) top 1001 to 10.000, iv) top 10.001 to 50.000, and v) books beyond top 50.000. Using these groups, we explore how many books we need to account for a given proportion of the drift over time, and how consistent the contributions of individual books are.
@@
Figure 3: Contributions of individual books to drift dynamics. A-B, Distributions of relative contributions of individual books to local drift (A) and global drift (B) over time. C, Probability that a book with a given level of contribution to the local drift in a month will have a certain level of contribution to the local drift in the next month. Probabilities are averaged across all pairs of months. D, Monthly popularity for the 1,000 books with the highest contributions to the global drift in October 2024. Books are vertically ordered by the month where their popularity peaks.
We start by examining how much each group of books contributes to the drift in collective attention over time (Fig. 3). Surprisingly, we find that the 100 books with the most significant shifts in their popularity consistently account for only around 10% of both local and global drift (Fig. 3A-B). This means that, even if it were possible to identify these books and predict changes in their popularity, 90% of the total shift in collective attention would remain unexplained without also accounting for less influential books. For local drift, the top 10,000 books that contribute most to drift account for less than 50% of the total change in collective attention. When it comes to seasonal variation in local drift (Fig. 3A), we observe a slight increases over time in the relative contributions of the top 101-1,000 highest contributing books, whereas the top 100 most contributing books do not increase their contributions. For global drift, we find that the relative contributions of the top 1,000 books decline over time (Fig. 3B). In other words, the 1,000 books with the largest changes in popularity are less important for explaining collective attention drift, the further we move from the reference month. Instead, shifts in the popularity of especially the top 1,001-10K most contributing books account for a larger proportion of global drift over time. Nonetheless, even when accounting for the top 10.000 most influential books, we explain less than 70% of global drift. Taken together, these results show that books in the long tail of the popularity distribution, in a collective fashion, are impactful in driving changes in collective attention.